# bibliography.bib

### v1

@@ 4.1 Motivation for SCM-based priors
Adding identifiability assumptions to the prior. A key question is whether we should actually impose the identifiability assumption $\kappa=0$ when constructing a prior. A different approach would be to also put a prior on $\kappa$, thus taking account the possibility of identifiability violations in the prior. Such an approach has been proposed by [Robertson.2025], where the authors construct a prior over many possible causal inference settings simultaneously. However, as we show in the following, this would make consistent Bayesian estimation of the causal query of interest impossible.
Lemma 4.1.
@@
See Appendix A. ∎
Lemma 4.1 has an important consequence: if our prior $\Pi$ puts positive probability mass on all possible combinations of $(\alpha,\beta,\gamma,\delta,\zeta,\eta,\theta,\kappa)$, the corresponding posterior $\Pi(\cdot\mid\mathcal{D}_{n})$ will even for $n\to\infty$ put positive probability mass on any $\zeta\in\mathbb{R}$, thus being completely non-informative about the causal target quantity. As a consequence, any Bayesian point estimator using such a prior (e.g., as in under the approach [Robertson.2025]) will be asymptotically biased.
In contrast, we present a different approach to circumvent the above problems: namely, we propose to construct PFN-priors that incorporate assumptions that allow for identifiability of the causal target quantity (e.g., setting $\kappa=0$ in the above example). As such, we follow established philosophy in causal inference that separates identifiability and estimation steps [Pearl.2009]: the identifiability step should be established by the practitioner using domain knowledge (e.g., establishing whether a certain variable is a valid instrument). Once identifiability has been established, we can use Bayesian modeling and PFN-based models for the estimation step.
Which noise variables to model? A key question that remains is what classes of SCMs we can use to specify priors for the causal inference setting $\mathcal{C}$ at hand. Indeed, the class of SCMs is non-unique: it turns out that it is not necessary to specify both noise variables $\epsilon_{A}$ and $\epsilon_{Y}$ as the next lemma shows.
Lemma 4.2.
@@
See Appendix A. ∎
Lemma 4.2 implies that it suffices to specify priors over SCMs without either treatment noise $\epsilon_{A}$ or outcome noise $\epsilon_{Y}$. However, if we remove both, there exist interventional distributions for which the prior will never put probability mass on the ground-truth causal query, rendering Bayesian inference inconsistent. In the following, we generalize this result to arbitrary SCMs and causal inference settings.
4.2 Formalizing SCM-based priors
In the previous example, we argued why sampling entire structural causal models (SCMs) can yield priors that are better aligned with causal queries than priors defined directly on observational distributions. We now provide a general formalism for such priors.
Definition 4.3 (SCMs [Pearl.2009]).
@@ 4.2 Formalizing SCM-based priors
Every SCM induces a unique directed acyclic graph (DAG), $\mathcal{G}^{\mathcal{S}}$ by defining mapping of the parents $pi(Z_{i})$ to $Z_{i}$ with directed edges. We distinguish two types of latent variables $U_{i}$ in $\mathcal{G}^{\mathcal{S}}$: $U_{i}$ is an unobserved confounder if it is the parent of both $A$ and $Y$, otherwise, we call it a noise variable.
Intuitively, an SCM is a simulator: we can draw latent variables $U\sim\mathbb{P}$ and pass them through structural functions $f$, resulting in an induced observational distribution $\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}}$ over $Z$. At the same time, we can modify the SCM by intevening on a variable via $do(A=a)$, i.e., fixing the variable and then sampling from the SCM mechanism. This induces a corresponding interventional distribution $\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}$. We call an SCM $\mathcal{S}$ compatible with a causal inference setting $\mathcal{C}$, if
$(\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}},\mathbb{P}_{\mathrm{int}}^{\mathcal{S}})\in\mathcal{P}_{\mathrm{obs}}\times\mathcal{P}_{\mathrm{int}}.$ (16)
Priors over SCMs.
We now define prior distribution over classes of SCMs that enable Bayesian causal inference.
Definition 4.4 ($\mathcal{C}$-SCM-Priors).
Sampling from $\Pi$ therefore amounts to sampling a random latent distribution $\mathbb{P}$ over $U$ as well as random functional assignments $f$. These can them ne used to intern sample an observational dataset $\mathcal{D}_{n}$, i.e., there is a well defined likelihood $\Pi(\mathcal{D}_{n}\mid\mathcal{S})$ induced by $\mathcal{S}$. As a consequence, we can define the posterior distribution over SCMs via
$\Pi(\mathcal{S}\mid\mathcal{D}_{n})\propto\Pi(\mathcal{D}_{n}\mid\mathcal{S})\Pi(\mathcal{S}),$ (17)
where $\propto$ denotes proportionality up to a normalization constant.
Well-specified priors.
The question is now how we should design our prior $\Pi$ such that the induced posterior $\Pi(\mathcal{S}\mid\mathcal{D}_{n})$ allows for valid Bayesian causal inference. We now define a key desirable property of such priors.
Definition 4.5 (Well-specified priors).
A prior $\Pi(\mathcal{S})$ is well-specified for $\mathcal{C}$ if, for every true pair $(\mathbb{P}_{\mathrm{obs}}^{\ast},\mathbb{P}_{\mathrm{int}}^{\ast})$ and every sample $\mathcal{D}_{n}\!\sim\!\mathbb{P}_{\mathrm{obs}}^{\ast}$,
Cluster-DAGs.
Because an SCM-prior induces a distribution over possibly many DAGs, we compress them into a shared structure.
Definition 4.6 (C-DAGs [Anand.2023]).
Given variables $\left(Z,U\right)$, a Cluster-DAG (C-DAG) is a DAG on clusters $C_{1},\dots,C_{k}$ which are disjoint subsets of $(Z,U)$.
Each $\mathcal{C}$-SCM-Prior induces a unique C-DAG via the following algorithm: (i) draw an edge $C_{i}\!\to\!C_{j}$ whenever any SCMs $\mathcal{S}$ with $\Pi(\mathcal{S})>0$ contains some arrow from any node of $C_{i}$ to any node of $C_{j}$ and no SCM $\mathcal{S}$ with $\Pi(\mathcal{S})>0$ contains some arrow from any node of $C_{j}$ to any node of $C_{i}$; (ii) merge $C_{i}$ and $C_{j}$ whenever both directions occur across SCMs $\mathcal{S}$ with $\Pi(\mathcal{S})>0$.
Every induced C-DAG inherits the well-specified proprty from its prior.
Definition 4.7 (Well-specified C-DAG).
A C-DAG $\mathcal{G}_{c}$ is well-specified for $\mathcal{C}$ if it is induced by some well-specified $\mathcal{C}$-SCM-Prior.
The following theorem provides a necissary condition on whether a C-DAG can be well-specified and generalizes our insights from Lemma 4.2.
Theorem 4.8.
If $\mathcal{G}_{c}$ is well-specified for $\mathcal{C}$ and there is an unobserved confounder $U$ between $A$ and $Y$ in $\mathcal{G}_{c}$, then, under the regularity assumptions A.1 in Appendix A, either $A$ or $Y$ must have a noise variable as parent in $\mathcal{G}_{c}$.
Proof.
See Appendix A. ∎
The theorem yields a practical design rule: leave at least one exogenous-noise variable on the treatment–outcome path. Doing so lets the prior span both confounded and unconfounded worlds, a prerequisite for consistent posterior causal inference.
4.3 Constructing SCM-based priors
Figure 1: C-Dags compatible with the three example causal inference settings. Yellow variables are observed, blue variables are unobserved, red variables are clusters.
We now propose a practical algorithm to construct $\mathcal{C}$-SCM-priors. We assume that we have access to a pair $(\mathcal{G}_{c},\mathcal{I})$, where $\mathcal{G}_{c}$ is a well-specified C-DAG for $\mathcal{C}$ and $\mathcal{I}$ is a set of constraints on SCMs $\mathcal{S}$ compatible with $\mathcal{G}_{c}$ ensuring that $\mathcal{S}$ is also compatible with $\mathcal{C}$.
Example 1: back-door adjustment. The observable variables are $(X,A,Y)$ together with noise variables. A compatible C-DAG is in Fig. 1 (left). The constraint set is
@@ 4.3 Constructing SCM-based priors
$\mathcal{I}(\mathcal{S})\;=\;\bigl{\{}\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}}(A=a\mid X=x)>0\bigr{\}},$ (19)
ensuring that all SCMs satisfy the positivity assumption.
Example 2: Front-door adjustment. Here, the obsesrved variables are $(X,A,M,Y)$ with noise variables and an unobserved confounder $U$ between $A$ and $Y$. A compatible C-DAG is in Fig. 1 (middle). The constraint set is
$\mathcal{I}(\mathcal{S})\;=\;\Bigl{\{}\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}}(A=a\mid X=x)>0,\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}}(M=m\mid X=x,A=a)>0\Bigr{\}},$ (20)
ensuring positivity for both treatments and mediators.
Example 3: Instrumental variables. The observed variables are $(X,Z,A,Y)$, augmented by noise variables and an unobserved confounder $U$ that is a joint parent of $A$ and $Y$ has no edge to the instrument $Z$; see the compatible C-DAG in Fig. 1 (right). The constraint set is
$\mathcal{I}(\mathcal{S})\;=\;\Bigl{\{}\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}}(Z=z\mid X=x)>0,f_{Y}^{\mathcal{S}}(X,A,U)=f(X,A)+g(X,U)\Bigr{\}}$ (21)
Overall algorithm. Given $(\mathcal{G}_{c},\mathcal{I})$ we propose to construct a prior distribution $\Pi$ over SCMs as follows: First, we order the clusters $(C_{1},\dots,C_{k})$ according to their hierarchy in the DAG (i.e., $C_{1}$ has no parents). Then, we iterate over each cluster $C_{i}$ as follows: if $C_{i}$ only contains latent variables, fix their distribution to a standard normal distribution via
$U^{(i)}\;\sim\;\mathcal{N}\!\bigl{(}0,\mathbf{I}\bigr{)}.$ (22)
If $C_{i}$ is a cluster of observed and latent variables, we assign a clustered BNN prior to $C_{i}$ (see below). If $C_{i}$ only contains observed variables, we assign a observational BNN prior.
Clustered BNN prior. For clusters that contain both observed and latent variables, we leverage a BNN-based prior inspired by TabPFN [Hollmann.2023]. This prior allows us to effectively sample potentially high-dimensional clusters of variables for which the internal causal structure is irrelevant to infer the causal query of interest. The prior is defined via
$g_{\theta}^{(i)}:\;\mathrm{pa}(C_{i})\;\longrightarrow\;\mathbb{R}^{r},\qquad\theta\sim\Pi_{C_{i}}\quad\text{subject to }g_{\theta}^{(i)}\text{ satisfying }\mathcal{I}(\mathcal{S}_{\theta}).$ (23)
We then sample random nodes from $g_{\theta}^{(i)}$ that coincide with observed nodes in $C_{i}$, while the remaining nodes serve as latent noise within the cluster. This corresponds to applying the approach taking in TabPFN [Hollmann.2023] to clusters $C_{i}$ in the C-DAG in which the causal structure does not matter for estimating our causal query.
Example 1: back-door adjustment. Here, the data distribution $\mathbb{P}$ can be seperated as follows: $(X,U_{X})\sim\mathbb{P}_{X}$ with $U_{X}$ denoting noise variables withing the cluster $X$, $U_{A}\sim\mathbb{P}_{U_{A}}$, $U_{Y}\sim\mathbb{P}_{U_{Y}}$, $A=f_{A}(X,U_{A})$, and $Y=f_{A}(X,A,U_{Y})$. Our algorithm proceeds as follows: $\mathbb{P}_{U_{A}}$ and $\mathbb{P}_{U_{Y}}$ are noise variables and are set to standard normal distributions. The cluster $(X,U_{X})$ contains both noise and observed variables, meaning that $\mathbb{P}_{X}$ is sampled from an clustered BNN prior. Finally, $A$ and $Y$ are observed variables meaning that $f_{A}$ and $f_{Y}$ are sampled from observational BNN priors.
Implementation details. (i) For each covariate cluster $C_{i}$ containing latent nodes, we sample a random MLP-style graph over $\mathrm{pa}\left(C_{i}\right)$ by drawing biases and edge-weights from $\Pi_{C_{i}}$ and then pruning edges at random to ensure acyclicity. We evaluate this graph with tanh activations and noise (from normal, uniform, Laplace, or logistic distribution) to produce continuous features, then apply randomized thresholds to discretize or binarize a subset, yielding mixed-type covariates via our unstructured BNN prior. (ii) For treatment (and outcome) clusters $C_{j}$ of purely observed nodes, we instantiate a second BNN $f_{\theta}^{(j)}$ over $\mathrm{pa}\left(C_{j}\right)$ (with $\theta\sim\Pi_{C_{j}}$ and the same acyclicity constraint). We forward-propagate the covariates through $f_{\theta}^{(j)}$ with injected noise to compute a scalar propensity score, then threshold to assign a binary treatment. We forward-propagate both covariates and treatment to obtain potential outcomes. The resulting treatment (and outcome) are sampled from our structured BNN prior.
5 CausalFM: Training
5.1 Training algorithm
Our goal is to train a PFN $q_{\theta}(Y\mid x)$ to approximate the conditional PPID
$\Pi_{\mathrm{int}}(Y\mid\mathcal{D}_{n},X=x)=\int\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}(Y\mid X=x)\,\Pi(\mathcal{S}\mid\mathcal{D}_{n})\,\mathop{}\!\mathrm{d}\mathcal{S}.$ (25)
Given an SCM prior $\Pi$ and a prior $\Pi_{N}$ over sample sizes, we propose the following modified PFN loss
Importantly, the dataset $\mathcal{D}$ is sampled from the observational distribution, while the pair $(X,Y)$ is sampled from the interventional distribution induced by a random SCM. This ensures that the PFN will aim to predict the interventional outcome $Y$ based on data following the observational distribution. A similar loss has been proposed by [Bynum.2025], which, however, is only based on the mean-squared error instead of the negative log-likelihood and thus does not allow an interpretation for approximating the PPID in a Bayesian setting. In particular, modeling the entire PPID allows us not only to provide point estimators of our causal query, but also account for uncertainty.
In practice, we sample a sample size $N_{j}\sim\Pi_{N}$, an SCM $\mathcal{S}_{j}\sim\Pi$ and an observational dataset $\mathcal{D}_{N_{j}}^{j}\sim\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{j}}$ by sampling from the SCM. Then, we modify the SCM by performing the intervention of interest (e.g., $do(A=a)$) and sample test points $(x_{j},y_{j})\sim\mathbb{P}_{\mathrm{int}}^{\mathcal{S}_{j}}$ from the interventional SCM. The approximated PFN-loss is then
^$\hat{\mathcal{L}}(\theta)=\sum_{j}\bigl{[}-\log q_{\theta}(y_{j}\mid\mathcal{D}_{N_{j}}^{j},x_{j})\bigr{]}.$ (27)
Example: back-door adjustment.
Here, sample a sample size $N_{j}\sim\Pi_{N}$, an SCM $\mathcal{S}_{j}\sim\Pi$ from our constructed prior distribution $\Pi$ and an observational dataset $\mathcal{D}_{N_{j}}^{j}\sim\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{j}}$. Then, we perform two interventions $do(A=1)$ and $do(A=0)$ to obtain test points $(x_{j},y_{j}(1)-y_{j}(0))\sim\mathbb{P}_{\mathrm{int}}^{\mathcal{S}_{j}}$. The PFN loss becomes
^$\hat{\mathcal{L}}(\theta)=\sum_{j}\bigl{[}-\log q_{\theta}(y_{j}(1)-y_{j}(0)\mid\mathcal{D}_{N_{j}}^{j},x_{j})\bigr{]}.$ (28)
Model architecture and training details. Our model implementation of $q_{\theta}(Y\mid x)$ builds upon the TabPFN architecture [Hollmann.2023]. Each observation in the dataset is tokenized during the embedding process. We use separate encoders for covariates, treatment, and outcome. These embedded tokens are processed through a transformer where attention is applied between the observations. The model uses a task-specific decoder to predict CATE. The classification head consists of a two-layer MLP with GELU activation. The model is trained with cross-entropy loss with learning rate $1\mathrm{e}{-3}$, weight decay $1\mathrm{e}{-5}$), batch size $16$, and sequence length $1024$. Training proceeds for up to $100$ epochs with early stopping after $10$ epochs of no improvement on the validation loss. We generate $10k$ synthetic datasets for training. Training our CausalFM model on an NVIDIA A100 GPU takes approximately $10$ hours. The full implementation details are in Appendix B.
6 Experiments
6.1 Set up
Baselines. We consider a broad array of state-of-the-art methods for treatment effect estimation from the literature. Considering the rich methods in the literature, we use the following baselines: (1) S-learner [kunzel2019metalearners]: the S-learner is a model-agnostic learner that trains a single regression model by concatenating the covariate and the treatment as input; (2) T-learner [kunzel2019metalearners]: the T-learner is a model-agnostic learner that trains separate regression models for treated and control groups; (3) X-learner [kunzel2019metalearners]: builds upon the T-learner by first imputing individual treatment effects in each group and then fitting models to these pseudo-effects—often yielding more accurate estimates when treatment and control group sizes are imbalanced. (4) TARNet [shalit2017estimating]: using representation learning to extract features of covariates and train separate branches for treated and control groups with regularization.(5) DR-learner [kennedy2023towards]: generates pseudo-outcomes based on the doubly-robust AIPW estimator; (6) RA-learner [curth2021nonparametric]: uses a regression-adjusted pseudo-outcome in the second stage. Further implementation details are in Appendix B.1.
Evaluation metrics. We use the precision of estimating the heterogeneous effects (PEHE) criterion [curth2021nonparametric, hill2011bayesian] to evaluate model performance on the CATE estimation task.
Table 1: Results for benchmarking CATE evaluation performance across 10 datasets. Reported: mean $\pm$ standard deviation over five-fold train-test splits.
$\mathrm{D}_{1}$ $\mathrm{D}_{2}$ $\mathrm{D}_{3}$ $\mathrm{D}_{4}$ $\mathrm{D}_{5}$ $\mathrm{D}_{6}$ $\mathrm{D}_{7}$ $\mathrm{D}_{8}$ $\mathrm{D}_{9}$ $\mathrm{D}_{10}$
S-learner [kunzel2019metalearners] 0.265 $\pm$ 0.005 0.375 $\pm$ 0.006 0.438 $\pm$ 0.005 0.577 $\pm$ 0.005 0.562 $\pm$ 0.005 0.726 $\pm$ 0.004 0.537 $\pm$ 0.007 0.472 $\pm$ 0.005 0.652 $\pm$ 0.008 0.685 $\pm$ 0.004
T-learner [kunzel2019metalearners] 0.271 $\pm$ 0.008 0.385 $\pm$ 0.010 0.458 $\pm$ 0.008 0.602 $\pm$ 0.012 0.587 $\pm$ 0.006 0.728 $\pm$ 0.004 0.623 $\pm$ 0.008 0.535 $\pm$ 0.011 0.581 $\pm$ 0.006 0.626 $\pm$ 0.010
TARNet [shalit2017estimating] 0.347 $\pm$ 0.024 0.727 $\pm$ 0.023 0.536 $\pm$ 0.023 0.844 $\pm$ 0.026 1.203 $\pm$ 0.018 0.782 $\pm$ 0.020 1.379 $\pm$ 0.018 0.929 $\pm$ 0.021 0.959 $\pm$ 0.019 1.211 $\pm$ 0.022
DR-learner [kennedy2023towards] 0.351 $\pm$ 0.013 0.558 $\pm$ 0.016 0.599 $\pm$ 0.012 0.734 $\pm$ 0.019 0.703 $\pm$ 0.015 0.781$\pm$ 0.016 0.726 $\pm$ 0.016 0.612 $\pm$ 0.010 0.781 $\pm$ 0.018 0.799 $\pm$ 0.016
RA-learner [curth2021nonparametric] 0.315 $\pm$ 0.015 0.407 $\pm$ 0.014 0.459 $\pm$ 0.013 0.546 $\pm$ 0.016 0.620 $\pm$ 0.011 0.641 $\pm$ 0.018 0.608 $\pm$ 0.017 0.522 $\pm$ 0.013 0.602 $\pm$ 0.012 0.655 $\pm$ 0.017
X-learner [kunzel2019metalearners] 0.280 $\pm$ 0.128 0.393 $\pm$ 0.124 0.458 $\pm$ 0.126 0.562 $\pm$ 0.018 0.602 $\pm$ 0.124 0.642 $\pm$ 0.130 0.609 $\pm$ 0.130 0.506 $\pm$ 0.127 0.628 $\pm$ 0.126 0.646 $\pm$ 0.125
CausalFM (ours) 0.263 $\pm$ 0.002 0.333 $\pm$ 0.001 0.417 $\pm$ 0.001 0.541 $\pm$ 0.003 0.559 $\pm$ 0.001 0.641 $\pm$ 0.001 0.508 $\pm$ 0.001 0.426 $\pm$ 0.001 0.561 $\pm$ 0.001 0.620 $\pm$ 0.002
Lower = better (best in bold).
Table 2: Results showing the CATE evaluation performance on the Jobs.
PEHE
S-learner [kunzel2019metalearners] 0.598 $\pm$ 0.004
T-learner [kunzel2019metalearners] 0.618 $\pm$ 0.008
TARNet [shalit2017estimating] 1.305 $\pm$ 0.025
DR-learner [kennedy2023towards] 0.663 $\pm$ 0.016
RA-learner [curth2021nonparametric] 0.611 $\pm$ 0.014
X-learner [kunzel2019metalearners] 0.801 $\pm$ 0.129
CausalFM (ours) 0.581 $\pm$ 0.002
Lower = better (best in bold).
Synthetic dataset. Due to the fundamental problem of causal inference, the ground-truth CATE can not be observed in real-world data. We thus generate synthetic datasets to evaluate model performance following prior literature (e.g., [kunzel2019metalearners, shalit2017estimating, kennedy2023towards]). We benchmark our model on ten datasets with different data generation mechanisms. Further details are provided in the Appendix B.
Jobs dataset. The study by LaLonde [lalonde1986evaluating] is a widely used benchmark in the causal inference field. This dataset combines a randomized study based on the National Supported Work program with observational data [smith2005does]. The treatment is job training and the outcomes are income and employment status after training. The study includes 8 covariates, such as age and education, as well as previous earnings. We use the Jobs dataset constructed by [shalit2017estimating], where the goal is to predict unemployment.
6.2 Results
Table 2 shows the benchmarking results on the Jobs dataset. Table 1 shows the benchmarking results on 10 synthetic datasets. The experiments demonstrate that our CausalFM achieves competitive CATE estimation performance on these datasets, without the need to re-train the model.
7 Conclusion and Future Work
This paper introduces CausalFM, a framework for training foundation models for causal inference using prior-data fitted networks (PFNs). While we instantiate the approach for estimating CATEs, this represents only a first step. The next steps are instantiations to broader causal inference settings like instrumental variables and front-door adjustment.
We are currently expanding our work, and, in the next step, we will instantiate other versions of CausalFM to broader causal inference settings as described above. In addition, we plan to scale CausalFM to larger and more diverse datasets to demonstrate the generalizability and practical applicability of our methods. Thereby, our aim to establish robust foundation models for a wide range of causal inference settings.
8 Acknowledgments
This work has been supported by the German Federal Ministry of Education and Research (Grant: 01IS24082).
[…]
Appendix A Proofs
A.1 Linear IV
Proof of Lemma 4.1.
@@ A.1 Linear IV
Choose $\kappa\neq 0$ such that $|\kappa|<|\alpha^{*}|$. Set:
$\alpha$ $=\sqrt{(\alpha^{*})^{2}-\kappa^{2}}$ (47)
@@
Since $\alpha\beta=\alpha\beta^{*}=\frac{\alpha}{\alpha^{*}}\alpha^{*}\beta^{*}$, we have:
$(\alpha\beta)^{2}=\frac{\alpha^{2}}{(\alpha^{*})^{2}}(\alpha^{*}\beta^{*})^{2}=\frac{(\alpha^{*})^{2}-\kappa^{2}}{(\alpha^{*})^{2}}(\alpha^{*}\beta^{*})^{2}$
Therefore:
$\kappa\beta+\gamma=\pm\sqrt{(\gamma^{*})^{2}+\frac{\kappa^{2}}{(\alpha^{*})^{2}}(\alpha^{*}\beta^{*})^{2}}$
From equations (44) and (45), we can solve for $\eta$:
$\eta=\frac{\zeta^{*}(\alpha^{*})^{2}\beta^{*}-\zeta(\alpha^{2}\beta+\kappa(\kappa\beta+\gamma))}{\kappa}$
Finally, $\theta$ is determined from equation (46).
@@
Proof of Lemma 4.2..
Let $\mathcal{S}=(\alpha,\beta,\gamma,\delta,\zeta,\eta,\theta)$ be any SCM from the linear IV class in Eq. (14). Then the following coefficients are identified via obserational data alone:
$\alpha=\sqrt{\mathrm{Var}(Z)},\beta=\mathbb{E}[Y\mid Z=1],\zeta=\frac{\zeta\beta}{\beta}=\frac{\mathbb{E}[\zeta(\beta+\delta\epsilon_{A})]-\mathbb{E}[\zeta(\delta\epsilon_{A})]}{\beta}=\frac{\mathbb{E}[Y\mid Z=1]-\mathbb{E}[Y\mid Z=0]}{\mathbb{E}[A|Z=1]-\mathbb{E}[A|Z=0]}$ (51)
@@
$\delta^{2}+\gamma^{2}=\mathrm{Var}(A\mid Z),\eta^{2}+\theta^{2}=\mathrm{Var}(Y\mid A)$ (52)
@@
Due to Eq. (53) it must holds that $\frac{\eta_{1}\gamma_{1}}{\gamma_{1}^{2}+(\beta_{1}\alpha_{1})^{2}}=\frac{\eta^{\ast}\gamma^{\ast}}{{\delta^{\ast}}^{2}+{\gamma^{\ast}}^{2}+(\beta^{\ast}\alpha^{\ast})^{2}}$, which implies that
@@
Due to Eq. (53) it must holds that $\frac{\eta\gamma}{\gamma^{2}+\delta^{2}+(\beta\alpha)^{2}}=\frac{\eta\gamma}{{\delta^{\ast}}^{2}+{\gamma^{\ast}}^{2}+(\beta^{\ast}\alpha^{\ast})^{2}}$ which implies that

### v2

@@ Appendix B Example for SCM-priors
Adding identifiability assumptions to the prior. A key question is whether we should actually impose the identifiability assumption $\kappa=0$ when constructing a prior. A different approach would be to also put a prior on $\kappa$, thus taking account the possibility of identifiability violations in the prior. Such an approach has been proposed by (Robertson et al., 2025), where the authors construct a prior over many possible causal inference settings simultaneously. However, as we show in the following, this would make consistent Bayesian estimation of the causal query of interest impossible, confirming Theorem 4.3.
Lemma B.1.
@@
See Appendix C. ∎
Lemma B.1 has an important consequence: if our prior $\Pi$ puts positive probability mass on all possible combinations of $(\alpha,\beta,\gamma,\delta,\zeta,\eta,\theta,\kappa)$, the corresponding posterior $\Pi(\cdot\mid\mathcal{D}_{n})$ will even for $n\to\infty$ put positive probability mass on any $\zeta\in\mathbb{R}$, thus being completely non-informative about the causal target quantity. As a consequence, any Bayesian point estimator using such a prior (e.g., as in under the approach (Robertson et al., 2025)) will be asymptotically biased.
In contrast, we present a different approach to circumvent the above problems: namely, we propose to construct PFN-priors that incorporate assumptions that allow for identifiability of the causal target quantity (e.g., setting $\kappa=0$ in the above example). As such, we follow established philosophy in causal inference that separates identifiability and estimation steps (Pearl, 2009): the identifiability step should be established by the practitioner using domain knowledge (e.g., establishing whether a certain variable is a valid instrument). Once identifiability has been established, we can use Bayesian modeling and PFN-based models for the estimation step.
Which noise variables to model? A key question that remains is what classes of SCMs we can use to specify priors for the causal inference setting $\mathcal{C}$ at hand. Indeed, the class of SCMs is non-unique: as suggested in the main paper, it is not necessary to specify both noise variables $\epsilon_{A}$ and $\epsilon_{Y}$.
Lemma B.2.
@@
See Appendix C. ∎
Lemma B.2 implies that it suffices to specify priors over SCMs without either treatment noise $\epsilon_{A}$ or outcome noise $\epsilon_{Y}$. However, if we remove both, there exist interventional distributions for which the prior will never put probability mass on the ground-truth causal query, rendering Bayesian inference inconsistent. In the following, we generalize this result to arbitrary SCMs and causal inference settings.
Appendix C Proofs
C.1 Proof of Theorem 4.3
Proof of Theorem 4.3.
@@ Appendix C Proofs
$Q(\mathbb{P}_{\mathrm{int}}^{\mathcal{W}})\;=\;\bar{Q}(\mathbb{P}_{\mathrm{obs}}^{\mathcal{W}})\;=\;\bar{Q}(\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{0}})\neq Q(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}_{0}}).$ (12)
Draw data $\mathcal{D}_{n}\sim\mathbb{P}_{\mathrm{obs}}^{\mathcal{W}}$. For every $\mathcal{S}\in\mathcal{E}(\mathbb{P}_{\mathrm{obs}}^{\mathcal{Z}})$, the observational likelihoods coincide for all $n$, so Bayes factors within $\mathcal{E}(\mathbb{P}_{\mathrm{obs}}^{\mathcal{Z}})$ are identically $1$. Consequently,
@@
$\Pi\!\bigl(\mathcal{S}\mid\mathcal{D}_{n}\bigr)\propto\Pi(\mathcal{S})\quad\text{for all }\ \mathcal{S}\in\mathcal{E}(\mathbb{P}_{\mathrm{obs}}^{\mathcal{Z}}),\ \text{all }n.$
We also know that $\Pi(\mathcal{S}|\mathcal{D}_{n})\to 0$ for $S\notin\mathcal{E}(\mathbb{P}_{\mathrm{obs}}^{\mathcal{Z}})$ and $n\to\infty$.
Hence,
$Q\left(\int\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}\,\Pi(\mathcal{S}\mid\mathcal{D}_{n})\,d\mathcal{S}\right)\;$ $\to\;\int_{\mathcal{E}(\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{0}})}Q(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}})\,\Pi(\mathcal{S})\,d\mathcal{S}$ (13)
$=w_{0}\,\int_{\mathcal{Z}}Q(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}})\,\Pi(\mathcal{S})\,d\mathcal{S}\;+\;(1-w_{0})\,Q\!\bigl(\mathbb{P}_{\mathrm{int}}^{\mathcal{W}}\bigr),$ (14)
$\neq Q\!\bigl(\mathbb{P}_{\mathrm{int}}^{\mathcal{W}}\bigr),$ (15)
which shows that $\Pi$ is not well-specified for $\mathcal{C}$. ∎
C.2 Proof of Lemma B.1 (linear IV)
Proof of Lemma B.1.
@@
We choose $\kappa\neq 0$ such that $|\kappa|<|\alpha^{*}|$. We set
$\alpha$ $=\sqrt{(\alpha^{*})^{2}-\kappa^{2}},$ (34)
@@
Since $\alpha\beta=\alpha\beta^{*}=\frac{\alpha}{\alpha^{*}}\alpha^{*}\beta^{*}$, we have .
$\alpha\beta)^{2}=\frac{\alpha^{2}}{(\alpha^{*})^{2}}(\alpha^{*}\beta^{*})^{2}=\frac{(\alpha^{*})^{2}-\kappa^{2}}{(\alpha^{*})^{2}}(\alpha^{*}\beta^{*})^{2}.$ (38)
Therefore, we have
$\kappa\beta+\gamma=\pm\sqrt{(\gamma^{*})^{2}+\frac{\kappa^{2}}{(\alpha^{*})^{2}}(\alpha^{*}\beta^{*})^{2}}.$ (39)
From Eq. (31) and Eq. (32), we can solve for $\eta$ via
$\eta=\frac{\zeta^{*}(\alpha^{*})^{2}\beta^{*}-\zeta(\alpha^{2}\beta+\kappa(\kappa\beta+\gamma))}{\kappa}.$ (40)
Finally, $\theta$ is determined from Eq. (33).
@@
Proof of Lemma B.2..
Let $\mathcal{S}=(\alpha,\beta,\gamma,\delta,\zeta,\eta,\theta)$ be any SCM from the linear IV class in Eq. (10). Then, the following coefficients are identified via observational data alone:
$\alpha$ $=\sqrt{\mathrm{Var}(Z)},$ (41a)
$\beta$ $=\mathbb{E}[Y\mid Z=1],$ (41b)
$\zeta$ $=\frac{\zeta\beta}{\beta}=\frac{\mathbb{E}[\zeta(\beta+\delta\epsilon_{A})]-\mathbb{E}[\zeta(\delta\epsilon_{A})]}{\beta}=\frac{\mathbb{E}[Y\mid Z=1]-\mathbb{E}[Y\mid Z=0]}{\mathbb{E}[A|Z=1]-\mathbb{E}[A|Z=0]}$ (41c)
@@
$\delta^{2}+\gamma^{2}=\mathrm{Var}(A\mid Z),\qquad\text{and}\qquad\eta^{2}+\theta^{2}=\mathrm{Var}(Y\mid A)$ (42)
@@
Due to Eq. (43), it must holds that
$\frac{\eta_{1}\gamma_{1}}{\gamma_{1}^{2}+(\beta_{1}\alpha_{1})^{2}}=\frac{\eta^{\ast}\gamma^{\ast}}{{\delta^{\ast}}^{2}+{\gamma^{\ast}}^{2}+(\beta^{\ast}\alpha^{\ast})^{2}},$ (47)
which implies that
@@
Due to Eq. (43), it must holds that $\frac{\eta\gamma}{\gamma^{2}+\delta^{2}+(\beta\alpha)^{2}}=\frac{\eta\gamma}{{\delta^{\ast}}^{2}+{\gamma^{\ast}}^{2}+(\beta^{\ast}\alpha^{\ast})^{2}}$ which implies that